# Visibility-Aware Keypoint Localization for 6DoF Object Pose Estimation

### v1

@@ 4.4 Ablation Study on LM Dataset
Ours ($N^{\prime}=256$) 48.6 85.9 97.0 77.2
We report ablation studies on LM [11] in Tab. 3. We train a unified pose estimator for all 13 objects. We first train the layers generating low-level quantities for 50k steps with learning rate of 2e-4, then we train all network layers for 100k steps, and finally reduce the learning rate to 1e-4 and train 20k steps.
Comparison with State of the Art. As shown in Tab. 3, our method significantly improves the performance w.r.t. ADD(-S) with threshold $2\%$ and $5\%$ (denoted as 0.02d and 0.05d). This demonstrates that our visibility-aware framework can greatly increase the ratio of qualified poses w.r.t. a strict threshold. Our method achieves comparable results w.r.t. ADD(-S) with threshold $10\%$ (denoted as 0.1d). The average of the metrics is the best among all methods.
Effectiveness of Visibility-Aware Keypoint Selection. We report the performance of using a fixed set of $N$ evenly distributed keypoints in localization step in Tab. 3. The results (denoted as w/o Selection) with $N=256$ and $N=512$ degrade especially for ADD(-S) 0.02d, which clearly demonstrates the effectiveness of our visibility-aware keypoint selection scheme.
Effectiveness of Dual Visibility Estimation. Our methods select keypoints based on both external visibility $V_{\mathrm{ex}}$ and internal visibility $V_{\mathrm{in}}$. Since the external occlusion is mild in test data of LM dataset, self-occlusion is the major factor of invisibility. As shown in Tab. 3, selecting keypoints based on $V_{\mathrm{in}}$ (denoted as w/o $V_{\mathrm{ex}}$) can significantly boost the performance compared with w/o Selection. Incorporating $V_{\mathrm{ex}}$ further improves the performance.
Effectiveness of Positional Encoding. In Tab. 3, we also present the result without positional encoding (denoted as w/o P. E.). Since the selected keypoints are dynamic w.r.t. input images, we find that adding positional encoding greatly improves the performance.
Effectiveness of Two-stage Training. In Tab. 3, we also report the performance without two-stage training (denoted as w/o Two-stage). The overall performance degrades without two-stage training, since the induced subgraph $G_{\mathrm{vis}}$ is not stable when the visibility estimator does not converge.
Number of Selected Keypoints. We show the results with different number of selected keypoints ($N^{\prime}$) in Tab. 3. The performance improves when $N^{\prime}$ is gradually increased towards 256, and remains almost the same when $N^{\prime}$ is increased to 320 with more computational cost. We use $N^{\prime}=256$ in practice.
4.5 Comparison to State of the Art
Table 4: Comparison on LM-O. We report the Average Recall (%) of ADD(-S) with three thresholds: 0.02d, 0.05d, and 0.1d. We highlight the best result and the second best result in red and blue, respectively.
Method GDR [64] Zebra [54] LC [33] Checker [31] Ours
@@ 4.5 Comparison to State of the Art
Following [64, 54, 31, 33], we train a single pose estimator for each object on LM-O and YCB-V datasets.
Experiments on LM-O. For the two-stage training procedure, we set the first stage as 50k steps and the second stage as 700k steps. We report the average recall of ADD(-S) metric with three thresholds (0.02d, 0.05d, and 0.1d) in Tab. 4, and provide the detailed results of the 8 objects in the Supplementary Material. As shown in Tab. 4, our visibility-aware keypoint localization scheme clearly boosts our base CheckerPose [31] on all three thresholds. Our method also greatly surpasses previous methods w.r.t. 0.05d threshold and the mean of three thresholds, and achieves comparable performance w.r.t. other thresholds.
Table 5: Comparison on YCB-V. We report the average ADD(-S), and AUC of ADD-S and ADD(-S). AUC values computed without and with 11-point interpolation are denoted as “w/o" and “w/ IT", respectively. “–" denotes unavailable results.
@@
Experiments on YCB-V. For the two-stage training procedure, we set the first stage as 50k steps and the second stage as 250k steps. We report the average values of ADD(-S) (0.1d) and AUC metrics in Tab. 5, and provide detailed results of the 21 objects in the Supplementary Material. Our method significantly improves the pose estimation performance w.r.t. the ADD(-S). Our method also achieves the best performance w.r.t. metrics based on AUC, which indicates that our method achieves the best accumulated performance across various thresholds.
4.6 Runtime Analysis
Table 6: Runtime Analysis. The unit of time is millisecond.
@@ 4.6 Runtime Analysis
In Tab. 6, we report the running speed of the methods that first establish dense correspondences and then use Progressive-X [2] as PnP solver. For an input $640\times 480$ RGB image, we evaluate the speed on a desktop with an Intel 2.30GHz CPU and an NVIDIA TITAN RTX GPU. The results show clear speed advantage of our method in addition to achieving state-of-the-art accuracies.
5 Conclusion
We propose a novel visibility-aware keypoint-based method, named VAPO, for instance-level 6DoF object pose estimation. By localizing the important keypoints guided by visibility, we efficiently invest computational resources to establish more reliable 3D-2D correspondences. From object-level annotations, we generate binary keypoint visibility labels as well as real-valued visibility-aware importance, for both asymmetric and symmetric objects. The extensive experiments on LM, LM-O and YCB-V datasets demonstrate that our method significantly improves both keypoint localization and object pose estimation.

### v5

@@ IV-B Ablation Study on LM Dataset
Ours ($k=15$) 48.0 86.1 97.1 77.1
Ours ($c=0.9$) 47.6 85.5 97.0 76.7
Ours ($c=0.85$) 48.6 85.9 97.0 77.2
Ours ($c=0.8$) 49.2 86.1 97.0 77.4
TABLE II: Comparison on LM-O. We report the Average Recall (%) of ADD(-S) with three thresholds: 0.02d, 0.05d, and 0.1d. We highlight the best result and the second best result in red and blue, respectively. “–” denotes unavailable results due to the absence of models for evaluation with the specified threshold.
Method GDR Zebra LC Checker 6D-Diff Ours
[18] [15] [45] [16] [28]
@@
TABLE III: Comparison on YCB-V. We report the average ADD(-S), and AUC of ADD-S and ADD(-S). AUC values computed without and with 11-point interpolation are denoted as “w/o” and “w/ IT”, respectively. We highlight the best result and the second best result in red and blue, respectively. “–” denotes unavailable results.
@@
TABLE IV: Reconstruction error on LM. We use Chamfer distance to evaluate single-view reconstruction and our proposed combination (Section III-E). All results are percentage of the object’s diameter for scale-invariant comparison.
View 1 2 3 4 5 6 7 8 Ours
ape 17.0 17.7 18.4 19.5 26.6 10.0 24.8 20.3 9.7
benchv. 21.9 17.3 14.4 12.2 20.0 6.8 20.2 13.9 10.9
camera 13.5 14.9 18.4 12.8 10.8 13.4 19.0 17.6 11.5
can 13.4 19.9 8.1 12.2 11.8 14.9 8.6 15.7 9.6
cat 17.3 21.6 9.5 13.9 15.5 17.9 15.9 8.7 8.2
driller 27.0 21.5 21.4 13.0 17.7 24.6 26.7 10.9 12.5
duck 18.4 14.4 15.7 10.0 8.4 11.3 20.8 13.5 10.1
eggbox 8.7 13.2 8.7 11.6 12.5 15.6 15.6 11.3 7.5
glue 19.9 14.4 9.8 8.8 9.5 13.8 13.7 14.7 10.1
holep. 21.4 12.6 19.5 17.2 15.5 19.1 13.8 19.5 9.5
iron 13.3 18.4 12.0 23.4 13.5 13.9 20.6 15.7 9.9
lamp 16.5 21.5 11.9 9.0 15.7 8.3 22.6 11.6 9.7
phone 24.8 26.8 7.7 18.0 20.8 15.0 24.0 14.6 14.4
TABLE V: CAD-Free Object Pose Estimation on LM. We report ADD(-S) (0.1d) and highlight the best result and the second best result in red and blue, respectively.
Method RLLG [29] One [30] One++ [31] GS [32] Ours
ape 52.9 11.8 31.2 71.0 83.4
benchv. 96.5 92.6 97.3 99.8 98.9
camera 87.8 88.1 88.0 98.2 95.9
can 86.8 77.2 89.8 97.7 99.4
cat 67.3 47.9 70.4 86.7 94.7
driller 88.7 74.5 92.5 96.2 96.7
duck 54.7 34.2 42.3 77.2 82.4
eggbox 94.7 71.3 99.7 99.6 99.6
glue 91.9 37.5 48.0 98.4 99.4
holep. 75.4 54.9 69.7 87.4 90.9
iron 94.5 89.2 97.4 99.2 95.3
lamp 96.6 87.6 97.8 98.9 99.1
phone 89.2 60.6 76.0 85.0 93.7
mean 82.9 63.6 76.9 92.0 94.6
Effectiveness of Visibility-Aware Keypoint Selection. We report the performance of using a fixed set of $N$ evenly distributed keypoints in localization step in Table I. The results (denoted as w/o Selection) with $N=256$ and $N=512$ degrade especially for ADD(-S) 0.02d, which clearly demonstrates the effectiveness of our visibility-aware keypoint selection scheme.
Effectiveness of Dual Visibility Estimation. Our methods select keypoints based on external visibility $V_{\mathrm{ex}}$ and internal visibility $V_{\mathrm{in}}$. Since the external occlusion is mild in LM, selecting keypoints based on $V_{\mathrm{in}}$ (denoted as w/o $V_{\mathrm{ex}}$) can significantly boost the performance compared with w/o Selection. Incorporating $V_{\mathrm{ex}}$ further improves the performance.
Effectiveness of Positional Encoding. Since the selected keypoints are dynamic w.r.t. input images, we find that adding positional encoding (denoted as w/o P. E.) greatly improves the performance.
Effectiveness of Two-stage Training. In Table I, we also report the performance without two-stage training (denoted as w/o Two-stage). The overall performance degrades without two-stage training, since the induced subgraph $G_{\mathrm{vis}}$ is not stable when the visibility estimator does not converge.
Number of Selected Keypoints. We show the results with different number of selected keypoints ($N^{\prime}$) in Table I. The performance improves when $N^{\prime}$ is gradually increased towards 256, and remains almost the same when $N^{\prime}$ is increased to 320 with more computational cost.
Choice of Hyperparameters. Due to computational constraints, we use $k=20$ to construct the $k$-NN graph. We also evaluate $k=15$, which achieves comparable results as shown in Table I. For the damping factor $c$, in addition to the commonly used value of 0.85, we also evaluate $c=0.9$ and $c=0.8$ to assess sensitivity in Table I. The ADD(-S) (0.1d) metric remains unchanged for different $c$, while ADD(-S) (0.02d) and ADD(-S) (0.05d) improve as $c$ decreases.
(a) Input Image
(b) GDR-Net [18]
(c) ZebraPose [15]
(d) Checker [16]
(e) Ours
(f) Keypoints
Figure 3: Qualitative results. In the input image (a), the red bowl is flipped and severely occluded. In (b)-(e), we render the bowl based on the predictions of different methods. Background pixels are changed to grayscale values for better visualization. In (f) we visualize the localization results of the selected keypoints, generated by CheckerPose (blue) and our method (green).
IV-C Comparison to State of the Art
Experiments on LM. As shown in Table I, our method significantly improves the performance w.r.t. ADD(-S) with threshold $2\%$ and $5\%$ (denoted as 0.02d and 0.05d). This demonstrates that our visibility-aware framework can greatly increase the ratio of qualified poses w.r.t. a strict threshold. Our method achieves comparable results w.r.t. ADD(-S) with threshold $10\%$ (denoted as 0.1d). The average of the metrics is the best among all methods.
Experiments on LM-O. We report the average recall of ADD(-S) metric with three thresholds (0.02d, 0.05d, and 0.1d) in Table II. As shown in Table II, our visibility-aware keypoint localization scheme clearly boosts our base CheckerPose [16] on all three thresholds. Our method also greatly surpasses previous methods w.r.t. 0.05d threshold and the mean of three thresholds, and achieves comparable performance w.r.t. other thresholds.
@@ IV-C Comparison to State of the Art
Experiments on YCB-V. The real-world test set of YCB-V includes diverse objects in novel and cluttered scenes with severe occlusion, which can evaluate generalizability in realistic settings. We report the average values of ADD(-S) (0.1d) and AUC in Table III. Our method significantly improves the pose estimation performance w.r.t. the ADD(-S). Our method also achieves the best performance w.r.t.  AUC-based metrics, which indicates that our method achieves the best accumulated performance across various thresholds.
IV-D CAD-Free Object Pose Estimation on LM Dataset
TABLE VI: Runtime Analysis. We highlight the best result and the second best result in red and blue, respectively.
@@ IV-D CAD-Free Object Pose Estimation on LM Dataset
TABLE VII: Analysis on the inner visibility computed by Eq. 3 on LM objects. “Binary” denotes the accuracy of the binary visibility labels, while “PPR” denotes the accuracy of the keypoints with top PPR-based importance.
Object Binary PPR Object Binary PPR Object Binary PPR
ape 87.6 93.4 benchv. 66.0 82.8 camera 83.5 92.0
can 83.5 93.0 cat 80.6 88.0 driller 86.8 89.5
duck 89.0 91.9 eggbox 82.0 91.6 glue 94.4 96.9
holep. 60.4 67.3 iron 87.7 91.9 lamp 80.4 84.5
phone 83.2 91.4
We evaluate our method in the CAD-free setting on the LM dataset. As shown in Table IV, our strategy of combining visible parts from multiple reconstructions can effectively mitigate errors in individual reconstructions and reduce sensitivity to view selection. Therefore, we use the combined reconstruction instead of single-view reconstruction to train a unified pose estimator for all 13 objects. After training the low-level layers for 50k steps with learning rate of 2e-4, we train all network layers for 200k steps and then reduce the learning rate to 1e-4 and train 200k steps. As shown in Table V, our method significantly outperforms recent CAD-free baselines in terms of the ADD(-S) (0.1d) metric, which demonstrates the capability of our method to handle the more challenging CAD-free setting.
IV-E Discussion
Qualitative Results. In Figure 3, the predictions of previous methods result in incorrect orientations or physical intersections between the bowl and the sugar box. In contrast, our visibility-aware method estimate accurate orientation and translation (Figure 3 (e)). We also visualize the keypoint localization results in Figure 3 (f), where CheckerPose [16] generates dispersed results with significant outliers.
Runtime Analysis. In Table VI, we report the running speed of the methods that use Progressive-X [19] as a PnP solver, on a desktop with an Intel 2.30GHz CPU and an NVIDIA TITAN RTX GPU. The results show clear speed advantage of our method in addition to achieving state-of-the-art accuracies.
Non-Convex Objects. For computational efficiency, we approximate internal visibility by Eq. 3. In Table VII, we analyze the accuracy of the labels on LM, which contains 13 non-convex objects. For each object, we uniformly sample 8 views and use ray tracing to compute the ground truth binary visibility labels, which is computationally expensive. For most objects, the accuracy exceeds 80%. We also examine the accuracy of the keypoints selected by the PPR-based algorithm and observe that PPR effectively mitigates labeling error introduced by the approximation in Eq. 3.
V CONCLUSIONS AND FUTURE WORK
We propose a novel visibility-aware keypoint-based method, named VAPO, for instance-level 6DoF object pose estimation with and without 3D CAD models. From object-level annotations, we generate binary keypoint visibility labels as well as real-valued visibility-aware importance, for both asymmetric and symmetric objects. The extensive experiments on LM, LM-O and YCB-V datasets demonstrate that our method significantly improves object pose estimation. In future work, we will explore efficient ways to generate visibility labels when object-level annotations are unavailable. We will also incorporate our visibility-aware importance into zero-shot pose estimation to enhance the visual and geometric embeddings of novel CAD models.